OpenStudy (anonymous):

pls help... rationalize the denominator of 12/(3+sqrt{5}+2sqrt{2}

OpenStudy (anonymous):

\[multiply \top and bottom by (3-\sqrt{5-2\sqrt{2}}then multiply again by (1-\sqrt{10)}\]